OpenStudy (anonymous):

Write an equation and a solution for the problem. A health food store sells oatmeal for $3.60 per pound and bran flakes for $4.80 per pound. How many pounds of each should be used to get a mixture of 30 pounds that sells for $4.00 a pound?

OpenStudy (anonymous):

This is what I have so far 3.6x+4.8y=120 and x+y=30

OpenStudy (aaronq):

that's it, now just solve the equations. Take x+y=30 isolate y, y=30-x substitute into the other equation and solve

OpenStudy (anonymous):

So 3.6x+4.8y30-x=120?

OpenStudy (aaronq):

3.6x+4.8(30-x)=120

OpenStudy (anonymous):

3.6x+4.8(30-x)=120 3.6x+144+30x=120 33.6x+144=120 Is this good so far?

OpenStudy (aaronq):

yes

OpenStudy (anonymous):

then i subtract 144 from both sides?

OpenStudy (aaronq):

yeah or 120 from both sides

OpenStudy (anonymous):

so its 33.6x+24?

OpenStudy (anonymous):

wait

OpenStudy (anonymous):

-24=33.6x?

OpenStudy (aaronq):

oh you know what i just noticed, This 3.6x+144+30x=120 was supposed to be 3.6x+144-30x=120 so -26.4x-144=120

OpenStudy (aaronq):

wait no: -26.4x+144=120 i'm 100% sure this time lol

OpenStudy (anonymous):

now what?

OpenStudy (anonymous):

subtract 144 from both sides now

OpenStudy (anonymous):

to get -24=-26.4

OpenStudy (anonymous):

-24=-26.4x

OpenStudy (anonymous):

be back in a bit

OpenStudy (anonymous):

i'll ask someone else if your not there

OpenStudy (aaronq):

theres a mistake somewhere, hold on a second

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

type it in while im gone and i'll correct it

OpenStudy (aaronq):

okay, 3.6x+4.8(30-x)=120 3.6x+144-4.8x=120 144-1.2x=120 -1.2x=120-144=-24 \(x=\dfrac{24}{1.2}\)

OpenStudy (aaronq):

Then use the value of x in one of the original equations to find y

OpenStudy (anonymous):

Thank You!

OpenStudy (aaronq):

no problem !